The measure of an angle is 9 degrees more than twice its complement, find the measure of each angle.
Answer by checkley71(8403)

x+(x+9)=90
2x+9=90
2x=90-9
2x=81
x=81/2
x=40.5 for the smaller angle.
x+9-49.5
proof
40.5+49.5=90
90=90